Find the remaining trigonometric functions of θ if csc(θ) = 17/8 and cos(θ) < 0.

Answer by stanbon(57219)

Find the remaining trigonometric functions of θ if csc(θ) = 17/8 and cos(θ) < 0.
----
csc positive and cos negative , so theta in QII
-----------------------------------------------------
csc = r/y ; so r = 17 and y = 8
---
therefore x = sqrt[17^2-8^2] = sqrt[225] = 15
---
But x is negative in QII so x = -15
----
sin = y/r = 8/17
cos = x/r = -15/17
tan = y/x = -8/15
---
csc = r/y = 17/8
sec = r/x = -17/15
cot = x/y = -15/8
=====================
Cheers,
Stan H.
=====================